# Search for the charged lepton flavor violating decay $\psi(3686)\to e^{\pm}\mu^{\mp}$

### v1

@@ Abstract
The number of background events from $\psi(3686)$ decay is estimated by analyzing the $\psi(3686)$ inclusive and exclusive MC samples. The equivalent statistics of the inclusive MC sample is about the same as the data and the exclusive MC sample is ten times larger. The scaled number of background events from $\psi(3686)$ decay in the signal region is estimated to be 1.6.
The number of background events from the continuum process is estimated using $e^{+}e^{-}$ collision data samples at surrounding energy points, such as $\sqrt{s}=3.65,~{}3.682$ and $3.773~{}\rm GeV$. The c.m. energy and cross section of the continuum process have a rough inverse square relationship. Therefore, the scaled number of continuum background events at the $\psi(3686)$ c.m. energy can be obtained from the different data samples by
@@
where $N^{k}_{\rm cont}$ is the number of background events that survive in the signal region of the data sample with c.m. energy corresponding to index $k$, while $\mathscr{L}_{k}$ and $\mathscr{L}_{\psi(3686)}$ are the integrated luminosities of the data sample at c.m. energy $k$ and at the $\psi(3686)$ peak energy, respectively. $\sqrt{s_{k}}$ and $\sqrt{s_{\psi(3686)}}$ are the c.m. energies of different data samples. By weighting with errors, the scaled number of background events from the different data samples, $N^{s,k}_{\rm bkg2}$, is combined to give $N^{s}_{\rm bkg2}=4.6$. The expected sum of the background events from $\psi(3686)$ decay and the scaled continuum production is estimated to be 6.2.
The detection efficiency for the signal process is estimated using the signal MC sample, which is determined to be $(24.18\pm 0.16)\%$. The distribution of the signal MC sample inside and outside the signal region are shown in Fig 3(a). After applying all the selections to the data from $2367.0\times 10^{6}$ $\psi(3686)$ events, eight $\psi(3686)\to e^{\pm}\mu^{\mp}$ candidates are observed in the signal region, as shown in Fig. 3(b).
@@
The systematic uncertainty primarily arises from the tracking and PID of electrons and muons, the TOF difference requirement, the photon veto, the $|\Delta\theta|$ and $|\Delta\phi|$ requirement, the Bhabha event veto, the signal region requirement, the signal MC generation model, the statistics of the signal MC sample, and the number of $\psi(3686)$ events in data. The uncertainties from all sources are listed in Table 1. Control samples of $\psi(3686)(e^{+}e^{-})\to e^{+}e^{-}$ and $\psi(3686)(e^{+}e^{-})\to\mu^{+}\mu^{-}$ decays are used to estimate the systematic uncertainties from different selection requirements. The differences between MC and data for these two control samples give the uncertainties from tracking and PID selections. We combine these uncertainties by comparing weighted and unweighted signal efficiency. The systematic uncertainty of tracking and PID for the electron and muon is $0.5\%$. The uncertainties from other requirements are estimated by comparing the efficiencies in MC and data for the two control samples, where the larger efficiency difference is taken as the systematic uncertainty for each selection criterion. The uncertainties from the TOF difference requirement, photon veto, $\Delta\theta$ and $\Delta\phi$ requirements, Bhabha events veto selection, and signal region requirement are determined to be $0.2\%$, $1.1\%$, $3.8\%$, $3.3\%$, and $1.7\%$, respectively. The uncertainty from the signal MC model is estimated by comparing signal MC samples generated with two different models, which are the VLL model and a uniform phase-space model ref:evtgen , respectively. The uncertainty of the signal MC model is $10.0\%$. The uncertainty from signal MC statistics is determined to be $0.4\%$ according to the number of events in the signal MC sample. The uncertainty in the number of $\psi(3686)$ events from different years is quoted as $0.5\%$ ref:numpsip . Assuming no correlation between these sources, the total systematic uncertainty is $11.4\%$.
@@
Eight candidate events are observed in the signal region from the full dataset, while six background events are expected. Therefore, no obvious excess is observed, and an UL on the BF $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})$ is estimated using the profile likelihood method ref:likelihood . The likelihood function is defined as
@@
$\cdot\mathcal{G}(N_{\psi(3686)},N^{\rm data}_{\psi(3686)},\sigma_{N^{\rm data}_{\rm\psi(3686)}}),$
where the number of observed events, $N_{\rm obs}$, follows a Poisson distribution ($\mathcal{P}$) with an expected number equal to the sum of signal events ($N^{\rm data}_{\rm\psi(3686)}\cdot\mathcal{B}\cdot\epsilon_{\rm sig}$) and background events ($N_{\rm bkg1}+N_{\rm bkg2}$). Here, $N^{\rm data}_{\rm\psi(3686)}$ is the number of $\psi(3686)$ events in the data sample, $\mathcal{B}$ is the BF of $\psi(3686)\to e^{\pm}\mu^{\mp}$, $\epsilon_{\rm sig}$ is the signal detection efficiency, and $N_{\rm bkg1}$ and $N_{\rm bkg2}$ are the expected numbers of background events from $\psi(3686)$ decay and continuum production, respectively. The detection efficiency, $\epsilon_{\rm sig}$, follows a Gaussian distribution with a mean given by the efficiency, which is determined from the MC sample, $\epsilon^{\rm MC}_{\rm sig}$, and a standard deviation of $\epsilon^{\rm MC}_{\rm sig}\cdot\sigma^{\rm MC}_{\rm sig}$. Here, $\sigma^{\rm MC}_{\rm sig}$ represents the combined statistical and systematic uncertainty of the efficiency. The number of different background ($N^{\rm\psi(3686)-MC}_{\rm bkg1}$, $N^{k}_{\rm cont}$) follows a Poisson distribution with expected values of $N_{\rm bkg1}/f_{1}$ and $N_{\rm bkg2}/f_{2}$, respectively, where $f_{1}$ and $f_{2}$ are the scale factors for different background sources. The number of $\psi(3686)$ events, $N_{\psi(3686)}$, follows a Gaussian distribution with a mean of $N^{\rm data}_{\psi(3686)}$ and a standard deviation of $\sigma_{N^{\rm data}_{\psi(3686)}}$.
@@
The UL of $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})$ is estimated by scanning the likelihood function of the BF. The scaled likelihood distribution, normalized by the maximum likelihood value $L_{\rm max}$, is shown in Fig. 4. By integrating the likelihood function in the physical region where $\mathcal{B}\geq 0$, the UL on its BF is set to be $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})<1.4\times 10^{-8}$ at the $90\%$ C.L.
This result can be used to constrain the Wilson coefficients of effective operators in an effective field theory that accommodates CLFV Hazard:2016fnc . Combining the results of $J/\psi\to e^{\pm}\mu^{\mp}$ ref::jpsi2emu and $\psi(3686)\to e^{\pm}\mu^{\mp}$, we get the most stringent constraints on the Wilson coefficients of $|C_{DL}^{e\mu}/\Lambda^{2}|~{}(|C_{DR}^{e\mu}/\Lambda^{2}|)$, $|C_{VL}^{ce\mu}/\Lambda^{2}|~{}(|C_{VR}^{ce\mu}/\Lambda^{2}|)$ and $|C_{TL}^{ce\mu}/\Lambda^{2}|~{}(|C_{TL}^{ce\mu}/\Lambda^{2}|)$, which are set to be $1.1\times 10^{-4}~{}{\rm GeV^{-2}}$, $1.0\times 10^{-6}~{}{\rm GeV^{-2}}$ and $0.47~{}{\rm GeV^{-2}}$, respectively.
In summary, based on $(2367.0\pm 11.1)\times 10^{6}$ $\psi(3686)$ events collected by the BESIII detector at the BEPCII collider, no significant excess of $\psi(3686)\to e^{\pm}\mu^{\mp}$ events is observed in the dataset compared to the expected background. The UL on its BF is determined to be $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})<1.4\times 10^{-8}$ at the 90% C.L. This is the first result in the search for CLFV in $\psi(3686)$ decays, which is helpful to constrain the parameters of new physics models, and provides new input for the effective field theory of CLFV ref::psip2emu_significance .
Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B50G670107; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374

### v2

@@ 4 Background study
The number of background events from $\psi(3686)$ decay is estimated by analyzing the $\psi(3686)$ inclusive and exclusive MC samples. The equivalent statistics of the inclusive MC sample is about the same as the data and the exclusive MC sample is ten times larger. The scaled number of background events from $\psi(3686)$ decay in the signal region is estimated to be $1.6\pm 0.4$.
The number of background events from the continuum process is estimated using $e^{+}e^{-}$ collision data samples at surrounding energy points, such as $\sqrt{s}=3.65,~3.682$ and $3.773~\rm GeV$. The c.m. energy and cross section of the continuum process have a inverse square relationship. Therefore, the scaled number of continuum background events at the $\psi(3686)$ c.m. energy can be obtained from the different data samples by
@@
where $N^{k}_{\rm cont}$ is the number of background events that survive in the signal region of the data sample with c.m. energy corresponding to index $k$, while $\mathscr{L}_{k}$ and $\mathscr{L}_{\psi(3686)}$ are the integrated luminosities of the data sample at c.m. energy $k$ and at the $\psi(3686)$ peak energy, respectively. $\sqrt{s_{k}}$ and $\sqrt{s_{\psi(3686)}}$ are the c.m. energies of different data samples. By weighting with errors, the scaled number of background events from the different data samples is combined to be $N^{s}_{\rm bkg2}=\sum_{k}(N^{s,k}_{bkg2}/\sigma^{2}_{k})/\sum_{k}(1/\sigma^{2}_{k})=4.6\pm 1.4$, where $\sigma_{k}$ is the error of $N^{s,k}_{bkg2}$. The expected sum of the background events from $\psi(3686)$ decay and the scaled continuum production is estimated to be $6.2\pm 1.4$.
5 Systematic uncertainties
@@
The systematic uncertainty primarily arises from the tracking and PID of electrons and muons, the TOF difference requirement, the photon veto, the $|\Delta\theta|$ and $|\Delta\phi|$ requirement, the Bhabha event veto, the signal region requirement, the signal MC generation model, the statistics of the signal MC sample, and the number of $\psi(3686)$ events in data. The uncertainties from all sources are listed in Table 1. Control samples of $\psi(3686)(e^{+}e^{-})\to e^{+}e^{-}$ and $\psi(3686)(e^{+}e^{-})\to\mu^{+}\mu^{-}$ decays are used to estimate the systematic uncertainties from different selection requirements. The differences between MC and data for these two control samples give the uncertainties from tracking and PID selections. We combine these uncertainties by comparing weighted and unweighted signal efficiency. The weighted signal efficiency is defined as $\epsilon_{sig}^{w}=\sum_{i=1}^{n}\left(\prod_{r}\left(\epsilon_{r}^{\rm~Data}(P_{(t)},\cos\theta)/\epsilon_{r}^{\rm~MC}(P_{(t)},\cos\theta)\right)/n\right)$, where $\epsilon_{sig}^{w}$ is the weighted efficiency, $\epsilon_{r}^{\rm~Data}(P_{(t)},\cos\theta)$ and $\epsilon_{r}^{\rm~MC}(P_{(t)},\cos\theta)$ are the efficiency of control samples at different momentum and polar angle region, respectively, $r$ denotes different selection criteria, $n$ is the number of events in control samples. The systematic uncertainty of tracking and PID for the electron and muon is $0.5\%$. The uncertainties from other requirements are estimated by comparing the efficiencies in MC and data for the two control samples, where the larger efficiency difference is taken as the systematic uncertainty for each selection criterion. The uncertainties from the TOF difference requirement, photon veto, $\Delta\theta$ and $\Delta\phi$ requirements, Bhabha events veto selection, and signal region requirement are determined to be $0.2\%$, $1.1\%$, $8\%, 3.3$$3\%$, and $1.7\%$, respectively. The uncertainty from the signal MC model is estimated by comparing signal MC samples generated with two different models, which are the VLL model and a uniform phase-space model [46], respectively. The uncertainty of the signal MC model is $10.0\%$. The uncertainty from signal MC statistics is determined to be $0.4\%$ according to the number of events in the signal MC sample. The uncertainty in the number of $\psi(3686)$ events from different years is quoted as $0.5\%$ [33]. Assuming no correlation between these sources, the total systematic uncertainty is $11.4\%$.
@@
6 Results
Eight candidate events are observed in the signal region from the full dataset, while six background events are expected. Therefore, no obvious excess is observed, and an UL on the BF $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})$ is estimated using the profile likelihood method [54]. The likelihood function is defined as
@@ 6 Results
[…]
where the number of observed events, $N_{\rm obs}$, follows a Poisson distribution ($\mathcal{P}$) with an expected number equal to the sum of signal events ($N^{\rm data}_{\rm\psi(3686)}\cdot\mathcal{B}\cdot\epsilon_{\rm sig}$) and background events ($N_{\rm bkg1}+N_{\rm bkg2}$). Here, $N^{\rm data}_{\rm\psi(3686)}$ is the number of $\psi(3686)$ events in the data sample, $\mathcal{B}$ is the BF of $\psi(3686)\to e^{\pm}\mu^{\mp}$, $\epsilon_{\rm sig}$ is the signal detection efficiency, and $N_{\rm bkg1}$ and $N_{\rm bkg2}$ are the expected numbers of background events from $\psi(3686)$ decay and continuum production, respectively. The detection efficiency, $\epsilon_{\rm sig}$, follows a Gaussian distribution with a mean given by the efficiency, which is determined from the MC sample, $\epsilon^{\rm MC}_{\rm sig}$, and a standard deviation of $\epsilon^{\rm MC}_{\rm sig}\cdot\sigma^{\rm MC}_{\rm sig}$. Here, $\sigma^{\rm MC}_{\rm sig}$ represents the combined statistical and systematic uncertainty of the efficiency. The number of different background ($N^{\rm\psi(3686)-MC}_{\rm bkg1}$, $N^{k}_{\rm cont}$) follows a Poisson distribution with expected values of $N_{\rm bkg1}/f_{1}$ and $N_{\rm bkg2}/f_{2}$, respectively, where $f_{1}$ and $f_{2}$ are the scale factors for different background sources. The number of $\psi(3686)$ events, $N_{\psi(3686)}$, follows a Gaussian distribution with a mean of $N^{\rm data}_{\psi(3686)}$ and a standard deviation of $\sigma_{N^{\rm data}_{\psi(3686)}}$. The values of the parameters in Eq. (3) are listed in Table 2.
Table 2: The values of the parameters for the likelihood function in Eq. (3).
Parameter Value
$N_{\rm obs}$ 88
$N^{\rm data}_{\rm\psi(3686)}$ $2367.0\times 10^{6}$
$\sigma_{N^{\rm data}_{\rm\psi(3686)}}$ $11.1\times 10^{6}$
$\epsilon^{\rm MC}_{\rm sig}$ $24.18\%$
$\sigma^{\rm MC}_{\rm sig}$ $11.4\%$
$N^{\psi(3686)-{\rm MC}}_{\rm bkg1}$ $18$
$N^{3.773}_{\rm cont}$ $10$
$N^{3.682}_{\rm cont}$ 11
$N^{3.650}_{\rm cont}$ 0
$f_{1}$ 0.0894
$f^{3.773}_{2}$ 7.604
$f^{3.682}_{2}$ 8.797
$f^{3.650}_{2}$ 0.471
The UL of $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})$ is estimated by scanning the likelihood function of the BF. The scaled likelihood distribution, normalized by the maximum likelihood value $\mathcal{L}_{\rm max}$, is shown in Fig. 4. By integrating the likelihood function in the physical region where $\mathcal{B}\geq 0$, the UL on its BF is set to be $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})<1.4\times 10^{-8}$ at the $90\%$ C.L.
This result can be used to constrain the Wilson coefficients of effective operators in an effective field theory that accommodates CLFV [55]. Combining the results of $J/\psi\to e^{\pm}\mu^{\mp}$ [14] and $\psi(3686)\to e^{\pm}\mu^{\mp}$, we get the most stringent constraints on the Wilson coefficients of $|C_{DL}^{e\mu}/\Lambda^{2}|~(|C_{DR}^{e\mu}/\Lambda^{2}|)$, $|C_{VL}^{ce\mu}/\Lambda^{2}|~(|C_{VR}^{ce\mu}/\Lambda^{2}|)$ and $|C_{TL}^{ce\mu}/\Lambda^{2}|~(|C_{TL}^{ce\mu}/\Lambda^{2}|)$, which are set to be $1.77\times 10^{-4}~{\rm GeV^{-2}}$, $1.72\times 10^{-6}~{\rm GeV^{-2}}$ and $7.75\times 10^{-1}~{\rm GeV^{-2}}$, respectively.
@@
7 Summary
In summary, based on $(2367.0\pm 11.1)\times 10^{6}$ $\psi(3686)$ events collected by the BESIII detector at the BEPCII collider, no significant excess of $\psi(3686)\to e^{\pm}\mu^{\mp}$ events is observed in the dataset compared to the expected background. The UL on its BF is determined to be $\mathcal{B}(\psi(3686)\to e^{\pm}\mu^{\mp})<1.4\times 10^{-8}$ at the 90% C.L. This is the first result in the search for CLFV in $\psi(3686)$ decays, which is helpful to constrain the parameters of new physics models, and provides new input for the effective field theory of CLFV [56].
8 Acknowledgments
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; Agencia Nacional de Investigación y Desarrollo de Chile (ANID), Chile under Contract No. ANID PIA/APOYO AFB230003; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B50G670107; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374